### v1

# The impact of cosmic filaments on the abundance of satellite galaxies

Abstract
The impact of cosmic web environments on galaxy properties plays a critical role in understanding galaxy formation. Using the state-of-the-art cosmological simulation IllustrisTNG, we investigate how satellite galaxy abundance differs between filaments and the field, with filaments identified using the DisPerSE algorithm. When filaments are identified using galaxies as tracers, we find that, across all magnitude bins, central galaxies in filaments tend to host more satellite galaxies than their counterparts in the field, in qualitative agreement with observational results from the Sloan Digital Sky Survey. The average ratios between satellite luminosity functions in filaments and the field are $3.49$, $2.61$, and $1.90$ in the central galaxy $r$-band magnitude bins of $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20, respectively. We show that much of this excess can be attributed to the higher host halo masses of galaxies in filaments. After resampling central galaxies in both environments to match the halo mass distributions within each magnitude bin, the satellite abundance enhancement in filaments is reduced by up to $79\%$. Additionally, the choice of tracers used to identify filaments introduces a significant bias: when filaments are identified using the dark matter density field, the environmental difference in satellite abundance is reduced by more than $70\%$; after further resampling in both magnitude and halo mass, the difference is further suppressed by another $\sim 60$$95\%$. Our results highlight the importance of halo mass differences and tracer choice biases when interpreting and understanding the impact of environment on satellite galaxy properties.
[…]
software: Astropy (Astropy Collaboration et al., 2013, 2018), DisPerSE (T. Sousbie, 2011; T. Sousbie et al., 2011), Matplotlib (J. D. Hunter, 2007), Numpy (C. R. Harris et al., 2020), Scipy (P. Virtanen et al., 2020).
show]yxmeng@bao.ac.cn
1 Introduction
[…]
At the same time, galaxies and their halos are embedded in the large-scale cosmic web, composed of knots (nodes), filaments, sheets, and voids (J. R. Bond et al., 1996). Understanding how these large-scale environments influence galaxy properties is a fundamental question in galaxy formation (e.g., O. Hahn et al., 2007a, b; H. Mo et al., 2010). Among these components, filaments are particularly important: they contain a substantial fraction of the web’s mass (e.g., $\sim 50\%$, see M. Cautun et al., 2014), connect the densest regions, and funnel dark matter and baryons into knots, thereby influencing the evolution of the galaxies that reside within them.
[…]
By contrast, the impact of filaments on satellite galaxy systems is less explored, and current results remain debated. Compared to less dense environments (e.g., the field, including sheets and voids), halos of similar mass in denser environments (e.g., filaments) are expected to experience a higher halo merger rate (O. Fakhouri & C.-P. Ma, 2009), potentially leading to a greater abundance of satellites. This expectation is supported by observations from Q. Guo et al. (2015) (hereafter G15), who, using Sloan Digital Sky Survey (SDSS) data, find that the satellite luminosity function (LF) of galaxies in filaments is significantly higher, by a factor of $\sim 2$, than that of galaxies in non-filament environments. A recent study by F. Markos Hunde et al. (2025), based on dark matter-only cosmological simulations, further supports this picture, confirming that host halos in filaments typically contain more subhalos than those in voids. However, O. Metuki et al. (2015), utilizing the Galaxies-Intergalactic Medium Interaction Calculation (GIMIC) suite of cosmological hydrodynamical simulations (R. A. Crain et al., 2009), find that the abundance of satellites exhibits only a weak dependence on the web environment. In addition, they find that the number of subhalos per halo for a given mass shows an even weaker environmental dependence. Thus, the impact of filamentary environments on satellite abundance remains uncertain, and further observational and numerical studies – particularly with recent hydrodynamical simulations – are necessary.
In this study, we utilize one of the state-of-the-art cosmological hydrodynamical simulations, IllustrisTNG (F. Marinacci et al., 2018; J. P. Naiman et al., 2018; D. Nelson et al., 2018, 2019; A. Pillepich et al., 2018a, b; V. Springel et al., 2018), which allows us to investigate the impact of filaments on satellite abundance with direct insights into the physical origins of environmental effects. The paper is organized as follows. In Section 2, we describe the details of the IllustrisTNG simulations, the filament identification algorithm, and the definition of environment. In Section 3, we present the impact of filamentary environment on satellite abundance, including comparisons with observations, the impact of magnitude distributions, and the influence of halo mass distributions. The effects of tracer choice are investigated in Section 4. Finally, we summarize our findings and conclude in Section 5.
2 Methods
2.1 Simulation
Table 1: Number of central galaxies in bins of $r$-band absolute magnitude ($M_{r,{\rm cen}}$) and different cosmic web environments.
Bins $M_{r,{\rm cen}}$ range Total Knot Galaxy tracers Dark matter tracers
Filament Field Filament Field
$M_{r,{\rm cen}}\sim-23$ [23.5,22.5)[-23.5,-22.5) 64 39 23 2 24 1
$M_{r,{\rm cen}}\sim-22$ [22.5,21.5)[-22.5,-21.5) 455 19 235 201 328 108
$M_{r,{\rm cen}}\sim-21$ [21.5,20.5)[-21.5,-20.5) 1654 47 379 1228 756 851
$M_{r,{\rm cen}}\sim-20$ [20.5,19.5)[-20.5,-19.5) 2628 104 316 2208 760 1764
The analysis in this study is based on the IllustrisTNG simulations (F. Marinacci et al., 2018; J. P. Naiman et al., 2018; D. Nelson et al., 2018; A. Pillepich et al., 2018a; V. Springel et al., 2018), a series of cosmological, gravo-magneto-hydrodynamical simulations performed with the AREPO code (V. Springel, 2010). We use the $z=0$ snapshot from the TNG100-1 hydrodynamical run, which offers both high resolution and a large volume, enabling robust statistical analyses. This simulation includes $1820^{3}$ dark matter particles and an equal number of initial gas cells within a periodic box of size 110.7 cMpc (comoving Mpc). The adopted cosmology follows Planck Collaboration et al. (2016), with parameters $\Omega_{\Lambda,0}=0.6911$, $\Omega_{\rm m,0}=0.3089$, $\Omega_{\rm b,0}=0.0486$, $\sigma_{8}=0.8159$, $n_{\rm s}=0.9667$, and $h=0.6774$. The mean baryonic particle/cell mass is $1.4\times 10^{6}~M_{\odot}$, and the dark matter particle mass is $7.5\times 10^{6}~M_{\odot}$. In terms of spatial resolution, the softening length for collisionless particles (i.e., dark matter and star) at $z=0$ is 740 pc, while the minimum gas softening length is 185 pc.
Halos and subhalos in the simulation are identified using the friends-of-friends (FoF, M. Davis et al., 1985) and SUBFIND (V. Springel et al., 2001; K. Dolag et al., 2009) algorithms, respectively. We define the host halo mass as $M_{\mathrm{200c}}$, the total mass enclosed within the virial radius $R_{\mathrm{200c}}$, within which the mean density is 200 times the cosmic critical density. The galaxy sample is based on the SUBFIND subhalo catalogue in the TNG100-1 run. We exclude subhalos of non-cosmological origin (e.g., fragments or clumps formed through baryonic processes within existing galaxies, which are marked with SubhaloFlag=0 in the SUBFIND subhalo catalog; see section 5.2 in D. Nelson et al. 2019) as well as low-mass subhalos containing fewer than 100 star particles, which corresponds to satellite galaxies with stellar masses of $M_{\mathrm{star}}\sim 10^{8}~M_{\odot}$ or $r$-band magnitudes $M_{r}\sim-14.6$ (with a typical scatter of $\Delta M_{r}=0.6$). In this study, the stellar mass and $r$-band magnitude of each SUBFIND structure are computed using star particles within a 3D radial aperture of 30 physical kpc, consistent with the SDSS Petrosian aperture. To compute the satellite abundance of each FoF halo, we construct corresponding satellite samples. The central galaxy of each FoF halo is defined as the most bound SUBFIND subhalo in the catalogue, while satellite galaxies are defined as the remaining subhalos that meet the above criteria and reside within the virial radius of their host halo.
For a more direct comparison with the observational results in G15, we consider the SDSS $r$-band absolute magnitude of galaxies, $M_{r}$, in this study. We use $M_{r}$ data from D. Nelson et al. (2018), which are based on a model that accounts for dust attenuation due to the distribution of neutral gas and metals (specifically, ‘Model C’ in D. Nelson et al. 2018). We include all central galaxies with magnitudes in the range $23.5\leq M_{r,{\rm cen}}\leq-19.5$, and divide them into four magnitude bins. The bright-end limit of $M_{r,{\rm cen}}=-23.5$ is motivated by G15, while the faint-end limit of $M_{r,{\rm cen}}=-19.5$ is chosen to maximize the sample size within the range of reliable numerical resolution. The number of central galaxies in each bin is listed in the third column of Table 1.
2.2 Filament identification
Figure 1: Filaments in a $30$ Mpc-thick slice of the TNG100-1 simulation, identified by the DisPerSE algorithm using galaxies with stellar masses $M_{\rm star}\geq 10^{9}~M_{\odot}$ as tracers. The grayscale background shows the dark matter density field. White circles mark the $3\times R_{\mathrm{200c}}$ vicinities of galaxy clusters with $M_{\mathrm{200c}}\geq 10^{13.5}~M_{\odot}$, which are defined as the knot environment. Red dots indicate the centers of galaxies residing in knots, which are excluded from our analysis. Black solid line segments show the filament spines defined by DisPerSE. Orange dots mark the centers of galaxies located within cylinders of radius of $R_{\rm filament}=1$ Mpc around the filament spines (indicated by the orange shaded regions), and defined as filament galaxies. The remaining galaxies, shown as blue dots, are field galaxies. Visually, despite some minor discrepancies, the filament spines overall capture the massive filamentary structures in the underlying dark matter density field.
In this study, we identify filaments using the DisPerSE algorithm (T. Sousbie, 2011; T. Sousbie et al., 2011), a widely used scale-free filament finder applicable to both simulation and observational data (see e.g., N. I. Libeskind et al., 2018; A. Rost et al., 2020, for comparisons of different filament identification methods). Below, we briefly describe how DisPerSE works and refer readers to the original paper for theoretical details (T. Sousbie, 2011).
[…]
Following observations, we use galaxies as tracers for filament identification. Similar to D. Galárraga-Espinosa et al. (2024), we select galaxies according to observational limits, balancing the identification of real structures with the suppression of spurious ones. Specifically, we trace filaments using galaxies (including both central and satellite galaxies) with stellar masses $M_{\mathrm{star}}\geq 10^{9}~M_{\odot}$, apply smoothing on both the estimated density field and the identified filaments (each smoothed once), and select filaments with a persistence ratio above the threshold of $2\sigma$. Note that we have tested a range of persistence ratio thresholds (from $1\sigma$ to $6\sigma$), with and without smoothing procedures applied, and confirmed that the conclusions presented in the following sections are not sensitive to these choices.
The identified filament spines in a slice of the simulation box are shown as black solid lines in Figure 1. Overall, the filament spines capture the massive filamentary structures of the underlying dark matter density field (shown as the grayscale background), despite minor mismatches in smaller features. These discrepancies arise primarily from the use of a 2D projection and from galaxies being a sparse tracer of the density field. The impact of tracers will be examined in Section 4.
2.3 Definition of environment
After identifying the filamentary spines, we further classify central galaxies into ‘filament’ and ‘field’ environments. To do this, we first exclude the galaxies residing in clusters (i.e., the ‘knot’ environment), as these environments have distinct effects on galaxy properties compared to filaments and fields. Specifically, central galaxies located within $3R_{\mathrm{200c}}$ of clusters with $M_{\mathrm{200c}}\geq 10^{13.5}~M_{\odot}$ are classified as ‘knot’ environment galaxies (i.e., galaxies marked by red dots in Figure 1) and are excluded from our galaxy samples. The numbers of ‘knot’ galaxies in different $M_{r,{\rm cen}}$ bins are listed in the fourth column of Table 1.
To assign central galaxies to filaments, we adopt the conclusion from W. Wang et al. (2024) that the typical filament boundary lies at $\sim 1~\mathrm{Mpc}$, and identify a galaxy as part of a filament if its center lies within a cylinder of radius $R_{\rm filament}=1~\mathrm{Mpc}$ around the filament spine (see also the discussions in Q.-R. Yang et al., 2025). Central galaxies not assigned to either filaments or knots are classified as field galaxies. In this work, the environment is defined for each central galaxy and its host halo. Since our analysis concerns the satellites associated with their corresponding central galaxies, satellites are not classified separately. Instead, they are assigned the environment of the centrals. The distribution of galaxies used as tracers, along with their assigned cosmic web environments (orange for filament and blue for field), is shown in Figure 1. The numbers of central galaxies in each magnitude bin for the filament and field environments are given in the fifth and sixth columns of Table 1, respectively.
[…]
3 Satellite luminosity functions
3.1 Environmental dependence
Figure 2: Environmental dependence of satellite LFs. From left to right, each column shows the results for central galaxies in the bins of $M_{r,{\rm cen}}\sim-23$ (blue), 22-22 (red), 21-21 (green), and 20-20 (yellow). In the upper panel of each column, solid and dashed lines represent simulated satellite LFs for centrals in filaments and fields, respectively. Shaded regions indicate Poisson errors computed using Eq. (2). Filled squares and open circles show observational results from G15. Bottom panels display the ratio of the filament to field satellite LFs, with errors propagated accordingly. The satellite LF in the $M_{r,{\rm cen}}\sim-23$ bin for field galaxies is sparsely sampled due to the limited number of objects, and is therefore excluded from the following analysis. For the three remaining bins, central galaxies in filaments tend to host more satellites than their counterparts in the field, in qualitative agreement with the observational results of G15.
The satellite LF quantifies the abundance of satellite galaxies as a function of their luminosity, normalized to that of their central galaxy. In the top row of Figure 2, we show the average satellite LFs for central galaxies in filaments (solid lines) and fields (dashed lines), separated into different $M_{r,{\rm cen}}$ bins. The horizontal axis represents the magnitude difference between a satellite and its central galaxy, defined as $M_{\rm diff}=M_{r,{\rm sat}}-M_{r,{\rm cen}}$, which is related to the luminosity ratio between the satellite and its central galaxy. Note that $M_{\rm diff}$ is typically a positive value, as the central galaxy is by definition the brightest galaxy in a group. For satellites of a given central galaxy, a smaller (larger) $M_{\rm diff}$ corresponds to a brighter (fainter) satellite. The vertical axis shows the average number of satellites per central galaxy per unit $M_{\rm diff}$ bin, calculated as
$\phi_{\rm sat}(M_{\rm diff})=\frac{\sum_{i=1}^{N_{\rm cen}}N_{{\rm sat},i}(M_{\rm diff})}{N_{\rm cen}\Delta M_{\rm diff}},$ (1)
where $N_{{\rm sat},i}(M_{\rm diff})$ is the number of satellites associated with the $i$-th central galaxy whose magnitude differences fall within the bin $M_{\rm diff}-\Delta M_{\rm diff}/2,M_{\rm diff}+\Delta M_{\rm diff}/2)$, and $N_{\rm cen}$ is the total number of central galaxies considered. We also estimate the uncertainty in the satellite LF, shown as shaded regions, based on Poisson errors, i.e.,
$\delta\phi_{\rm sat}(M_{\rm diff})=\frac{\sqrt{\sum_{i=1}^{N_{\rm cen}}N_{{\rm sat},i}(M_{\rm diff})}}{N_{\rm cen}\Delta M_{\rm diff}}.$ (2)
To compare with observations, we overplot the observed satellite LFs from G15 in gray, using filled squares for filament galaxies and open circles for field galaxies from the SDSS. Note that G15 provide observational results only for the three brightest $M_{r,{\rm cen}}$ bins, and the $M_{r,{\rm cen}}\sim-21$ bin is affected by significant noise due to the small sample size.
As expected, moving from brighter to fainter central galaxies (from left to right panels), the abundance of satellites in the TNG100 simulations generally decreases in both filament and field environments. This trend is consistent with the observational results from G15.
In the brightest central galaxy bin ($M_{r,{\rm cen}}\sim-23$), there is only two central galaxy in our field sample. As a result, the corresponding satellite LF is sparsely sampled. Given the poor sampling of field central galaxies in this bin, we focus on the three fainter central galaxy bins in the following analysis to compare environmental differences more robustly.
For all three remaining bins in the simulation, we find that central galaxies in filaments tend to host more satellites than their counterparts in the field, in qualitative agreement with the trend in observational results of G15. This also echoes the similar environmental dependence of subhalo mass functions reported by F. Markos Hunde et al. (2025). In each $M_{r,{\rm cen}}$ bin, the environmental effect is more pronounced for brighter satellites (i.e., those with smaller $M_{\rm diff}$). When comparing across the $M_{r,{\rm cen}}$ bins, the difference between filament and field environments decreases toward fainter central galaxies. Specifically, in the $M_{r,{\rm cen}}\sim-22$ bin, the satellite abundance in filament is $\sim 2$66 times higher than in the field, with an average ratio of $3.49$. The ratio decreases to $\sim 2$44 (average $2.61$) in the $M_{r,{\rm cen}}\sim-21$ bin and $\sim 1$33 (average $1.90$) in the $M_{r,{\rm cen}}\sim-20$ bin. These average ratios are summarized in the third column of Table 2.
Overall, the environmental difference is quantitatively more pronounced in simulations compared to the G15 observations. This discrepancy could originate from several factors: (i) In simulations, central and satellite galaxies are very well-defined, while in observations, their identification can be affected by the uncertainties in galaxy redshifts and projection effects. Specifically, in G15, central galaxies are defined to have a characteristic radius, $R_{\rm inner}$, which is comparable to the virial radius of galaxies in the corresponding $M_{r,{\rm cen}}$ bin. Galaxies are selected as centrals if, within a projected distance of $2R_{\rm inner}$, all other galaxies are either more than half a magnitude fainter or have a spectroscopic redshift difference larger than 0.002 (or a photometric redshift difference larger than 2.5 times the photometric error). Satellite galaxies are more difficult to define, as they are fainter; thus, a statistical background subtraction technique is adopted to estimate their abundance. We refer interested readers to G15 for more details. The key message is that the definition of central and satellite galaxies in observations is more challenging, which could potentially affect the final results. (ii) The use of different filament identification methods could also contribute to the differences. G15 identify filament spines using the Bisous method (E. Tempel et al., 2014), an object point process model that detects and characterizes filamentary structures in the cosmic web by using a network of small, interacting cylindrical segments to trace filaments in galaxy distributions. Here, we use the DisPerSE finder, which is more widely used, open-source, and enables us to study the effects of different tracers. The differences in methods and parameter choices could also lead to quantitative discrepancies in the results.
In this work, we do not aim to fully reproduce the observations by strictly following the same procedures. The intriguing result is that the simulations exhibit a qualitatively similar environmental trend to that observed. This suggests that the trend is not tied to a particular choice of galaxy classification or filament identification, but instead emerges consistently, even though the methods employed here differ from those in G15. In the following subsections, we further examine the possible origin of this environmental difference.
Figure 3: Top: Distributions of central galaxy magnitudes ($M_{r,{\rm cen}}$). From left to right, we show the distributions of the original samples, the samples after resampling $M_{r,{\rm cen}}$, and the samples after resampling both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$. In each panel, as in other figures in this study, the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins are plotted in red, green, and yellow, respectively. Central galaxies in filaments and the field are distinguished using solid and dashed lines, respectively. Bottom: Same as the top panels, but showing the distributions of the host halo virial masses ($M_{\rm 200c}$) for central galaxies. After resampling, the filament and field samples exhibit similar distributions in both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$. See the main text for details of the resampling procedure.
Figure 4: Impact of magnitude and halo mass distributions on the environmental differences in satellite LFs. Left: The upper panel summarizes the satellite LFs for the $M_{r,{\rm cen}}\sim-22$ (red), 21-21 (green), and 20-20 (yellow) bins from Figure 2. Results for filament and field central galaxies are shown with solid and dashed lines, respectively. The lower panel shows the corresponding ratios between filament and field satellite LFs. Middle: Same as the left panels, but for galaxy samples resampled to have similar $M_{r,{\rm cen}}$ distributions. Right: Same as the middle panels, but for galaxy samples resampled to match both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$ distributions. After resampling $M_{r,{\rm cen}}$, the average ratios decrease by less than $17\%$. However, after additionally resampling $M_{\rm 200c}$, the average ratios drop by $78\%$, $79\%$, and $53\%$ for the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins, respectively, relative to the original unresampled samples. This indicates that differences in halo mass distributions between filament and field central galaxies contribute significantly to the environmental differences in satellite abundance, especially in the brighter central galaxy bins.
Table 2: Average satellite abundance ratios between filament and field central galaxies, $\phi_{\rm sat,filament}/\phi_{\rm sat,field}$. The values are calculated by averaging the ratios over different $M_{\rm diff}$ bins shown in Figures 4 and 6. For the resampled cases, the reductions of the ratios compared to the original cases are given in brackets, where the fractional reduction is computed as $1-\left(\frac{\phi_{\rm sat,filament}}{\phi_{\rm sat,field}}\big|_{\rm resampled}-1\right)\big/\left(\frac{\phi_{\rm sat,filament}}{\phi_{\rm sat,field}}\big|_{\rm original}-1\right)$. These results highlight that both halo mass and tracer choice significantly affect the measured environmental differences.
Filament tracers Bins $\phi_{\rm sat,filament}/\phi_{\rm sat,field}$
Original Resampling $M_{r,{\rm cen}}$ Resampling $M_{r,{\rm cen}}$ and $M_{\mathrm{200c}}$
Galaxies $M_{r,{\rm cen}}\sim-22$ $3.49$ $3.07~(17\%\downarrow)$ $1.55~(78\%\downarrow)$
$M_{r,{\rm cen}}\sim-21$ $2.61$ $2.54~(5\%\downarrow)$ $1.34~(79\%\downarrow)$
$M_{r,{\rm cen}}\sim-20$ $1.90$ $1.92~(2\%\uparrow)$ $1.43~(53\%\downarrow)$
Dark matter density $M_{r,{\rm cen}}\sim-22$ $1.67$ $1.34~(50\%\downarrow)$ $1.03~(95\%\downarrow)$
$M_{r,{\rm cen}}\sim-21$ $1.48$ $1.38~(20\%\downarrow)$ $1.02~(96\%\downarrow)$
[…]
3.2 Impact of magnitude distributions
In the previous analysis, we adopted a bin width of $\Delta M_{r,{\rm cen}}=1$ for central galaxies. A natural question is whether the observed environmental difference arises from the different distributions of $M_{r,{\rm cen}}$ within each bin between filament and field central galaxies. For example, if central galaxies in filaments have a higher fraction of brighter ones, they would naturally host more satellites.
To examine this, we first plot the histograms of central galaxy magnitudes in the upper-left panel of Figure 3 for the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins. Solid and dashed lines show the results for central galaxies in filaments and the field, respectively. Within each $M_{r,{\rm cen}}$ bin, central galaxies in fields tend to have a higher probability of being fainter, with the trend becoming more pronounced in the brighter $M_{r,{\rm cen}}$ bins. In contrast, central galaxies in filaments exhibit a relatively flat $M_{r,{\rm cen}}$ distribution in the 20-20 and 21-21 bins, while in the 22-22 bin, they have a slightly higher probability of being fainter, resembling their field counterparts. Overall, central galaxies in filaments do show a higher fraction of brighter members compared to those in the field, which can partially explain their higher satellite abundance.
To further quantify the impact of magnitude distribution on the environmental difference observed in Section 3.1, we resample the central galaxies based on their $M_{r,{\rm cen}}$, ensuring that the resampled samples have similar distributions in both environments. Specifically, for each pair of filament and field histograms, we divide them into 20 narrower bins. [^1] We have tested other choices for the number of narrow bins (e.g., 10, 40) and confirmed that the results remain consistent. A similar test is also performed for the resampling of $M_{\rm 200c}$ (Section 3.3). In each narrow bin, central galaxies from the environment with a higher number count are randomly resampled to match the number of galaxies in the other environment. The $M_{r,{\rm cen}}$ distributions of the resampled central galaxy samples are shown in the upper-middle panel of Figure 3. We can clearly see that, after resampling, the filament and field histograms are closely matched. We then re-examine the environmental differences in satellite LFs using these resampled central galaxies.
The satellite LFs computed from resampled central galaxies across different magnitude bins are plotted in the middle column of Figure 4. Compared to the original results shown in the left column, the resampled LFs exhibit larger Poisson errors due to reduced number counts. However, the ratio panels reveal that the environmental differences remain largely unchanged, with average ratios of $3.07$, $2.54$, and $1.92$ for the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins, respectively (see the fourth column in Table 2). Compared to the original ratios, the resampled values decrease by less than $17\%$. This suggests that the difference in magnitude distributions accounts for only a small part of the observed environmental effect, and that the environmental difference reported in Section 3.1 likely arises from other factors.
3.3 Impact of halo mass distributions
Another factor that can impact the abundance of satellite galaxies is the host halo mass, $M_{\rm 200c}$ (e.g., R. E. Angulo et al., 2009; M. Boylan-Kolchin et al., 2010; L. Gao et al., 2011; A. Rodríguez-Puebla et al., 2016). For example, if central galaxies in filaments reside more frequently in high-mass halos, they would naturally host more satellites. Unlike observations, hydrodynamical simulations provide precise measurements of host halo virial masses, allowing us to directly examine the impact of halo mass distributions.
In the bottom-left panel of Figure 3, we show histograms of the host halo masses of central galaxies across different environments and magnitude bins. Compared to the field, central galaxies in filaments are indeed more likely to reside in higher-mass halos, and this trend is especially pronounced in the brighter magnitude bins.
The bottom-middle panel presents the halo mass distributions after resampling the central galaxies to match their $M_{r,{\rm cen}}$ distributions. After this magnitude matching, the halo mass distributions in the two environments become somewhat closer, though notable differences remain. Specifically, in the $M_{r,{\rm cen}}\sim-20$ and 21-21 bins, filament galaxies tend to reside in more massive halos. In the 22-22 bin, the difference becomes more significant – filament galaxies are associated with more high-mass halos and fewer low-mass halos than their field counterparts. This indicates that even after matching the magnitude distributions, the halo mass distributions between filaments and the field remain substantially different, which could contribute to the environmental differences observed in the satellite LFs.
To quantify the impact of halo mass distributions, we follow the procedure described in Section 3.2. Starting from the resampled central galaxy samples with matched $M_{r,{\rm cen}}$ distributions, we further divide the halo mass distributions (shown in the bottom-middle panel of Figure 3) into 20 narrower bins. Within each narrow bin, we randomly resample central galaxies in the two environments to ensure equal numbers. This procedure yields central galaxy samples with very similar distributions in both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$, as shown in the right column of Figure 3.
With this newly resampled central galaxy samples, we again compute the satellite LFs and quantify the environmental differences. The results are shown in the right column of Figure 4. We find that the environmental differences are now substantially reduced. The average satellite abundance ratios decrease to $1.55$, $1.34$, and $1.43$ for the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins, respectively (see the fifth column of Table 2). Compared to the original samples, these values represent reductions of $78\%$, $79\%$, and $53\%$ , respectively. This suggests that the differences in halo mass distributions between filament and field central galaxies make a significant contribution to the environmental differences in satellite abundance observed in the original samples, especially in the brighter central galaxy bins.
3.4 Discussions
As shown in the previous subsections, the environmental differences in satellite abundances are partly driven by variations in both central galaxy magnitude and host halo mass distributions. Specifically, after resampling to match the $M_{r,{\rm cen}}$ distributions in the two environments, the differences in their satellite LFs are reduced by up to $17\%$. When the central galaxies are further resampled to match the $M_{\rm 200c}$ distributions, the environmental differences in satellite LFs decrease by up to $79\%$. This has key implications for interpreting observational results:
Our results suggest that difference in host halo mass distributions (i.e., filaments could host more massive halos compared to the field) can account for up to half of the observed environmental effects. However, measuring host halo masses precisely in observations is challenging.
The remaining environmental differences can be explained by two primary factors: the environmental dependence of halo merger rates (O. Fakhouri & C.-P. Ma, 2009) and the choice of tracers in filament identification. (i) Satellite galaxies originate from progenitor halos that have merged into a larger host and survived tidal disruption, so their abundance reflects the host halo’s merger history. Because merger rates correlate with large-scale density, and filaments are overdense compared to the field, halos in filaments are expected to experience more frequent mergers, naturally leading to a higher abundance of satellites. (ii) As we will show in the following section, the choice of tracer used to identify filaments also plays a significant role in the measured environmental differences in satellite abundance. In particular, using galaxies as tracers may enhance the environmental dependence of satellite LFs.
4 Tracer Effects
Figure 5: Similar to Figure 1, but for filaments identified using dark matter density as tracers. The spines of dark matter-traced filaments are shown as purple solid lines, while the galaxy-traced filament spines from Figure 1 are overlaid using black dotted lines for comparison. Filaments identified by these two tracers are qualitatively similar, both following the underlying filamentary matter distribution, but they differ in detailed structures. We investigate these differences and their impact in Section 4.
Figure 6: Same as Figure 4, but for the parallel analysis where filaments are identified using dark matter density as the tracer. Compared to Figure 4, where filaments are identified using galaxies, the environmental differences are significantly smaller here, suggesting that the choice of filament tracer also influences the strength of the environmental effect.
Figure 7: Number ratios between satellite-rich and satellite-poor central galaxies ($N_{\rm rich}/N_{\rm poor}$) as a function of total filament length normalized to the fiducial case ($L_{\rm tot,test}/L_{\rm tot,fiducial}$; a proxy for the persistence ratio threshold). Results from different central galaxy magnitude bins are shown from left to right. Solid lines represent dark matter-traced filaments, while dotted lines correspond to galaxy-traced filaments. Shaded regions indicate uncertainties estimated from Poisson errors. The vertical gray dashed line marks fiducial parameter we choose, with markers highlighting the corresponding ratios. Across all magnitude bins and filament lengths (spanning more than an order of magnitude by varying persistence thresholds), galaxy-traced filaments consistently show a preference for hosting satellite-rich central galaxies. In contrast, dark matter-traced filaments exhibit ratios close to unity (gray dashed horizontal line), indicating no significant preference.
4.1 Dark matter-traced filaments
It has been shown that the choice of tracers can affect the detailed properties of filaments identified by DisPerSE (see e.g., C. Laigle et al., 2018; D. Zakharova et al., 2023; Y. M. Bahe & P. Jablonka, 2025). Using cosmological simulations, these studies demonstrate that filament properties – such as spatial distributions, lengths, and connectivity – depend on whether dark matter particles or galaxies are used as tracers. This discrepancy arises because galaxies, distributed sparsely and discontinuously, are biased tracers of the underlying matter distribution, whereas dark matter more faithfully traces the total mass density field. In observations, we are limited to using galaxies as tracers, as was done in Section 3. In this section, we take advantage of simulations to explore the use of dark matter as tracers and investigate how the tracer choice affects the observed environmental dependence of satellite galaxy abundance.
For dark matter tracers, following Y. M. Bahe & P. Jablonka (2025), we use a smoothed density field instead of randomly sampled dark matter particles. This choice has minimal impact on the resulting filaments but yields more consistent identification with fewer artifacts at comparable computational expense. Specifically, we generate a gridded 3D density field with a cell size of 83.3 kpc, apply Gaussian smoothing with a scale of 500 kpc, and then downsample the field to a coarser grid with 250 kpc cells to reduce computational cost. Note that, since the numbers of tracer particles differ between different choices, it is usually not possible to use identical parameter sets in the DisPerSE finder. Following D. Zakharova et al. (2023), and to enable a more direct and fair comparison with galaxy-traced filaments, we adopt a higher persistence ratio threshold of $5.1\sigma$ (compared to $2\sigma$ for galaxy tracers), ensuring that the total filament lengths are matched between the two cases (i.e., $L_{\rm tot,fiducial}=7.84\times 10^{3}$ Mpc).
The resulting dark matter-traced filament spines are shown in Figure 5 as purple solid lines. For comparison, we also overplot the galaxy-traced filament spines using black dashed lines. The two filament networks appear qualitatively similar; however, the dark matter-traced filaments tend to be smoother and more closely follow the underlying dark matter distributions. Especially, the dark matter-traced filaments tend to extend more continuously, covering more low-density ridges that galaxy-traced filaments may miss.
4.2 Tracer effects on satellite luminosity functions
Once the dark matter-traced filament spines are identified, we follow the same environmental definitions outlined in Section 2.3 and apply the analysis described in Section 3, as done for the galaxy-traced case. The numbers of central galaxies classified into filament and field environments across different magnitude bins are given in the two rightmost columns of Table 1. Compared to the galaxy-traced case, a larger number of central galaxies are assigned to the filament environment, reflecting the fact that dark matter-traced filaments can capture more low-density ridges and thus include the galaxies surrounding these structures. The resulting environmental dependence of satellite LFs is summarized in Figure 6 and Table 2.
As shown in the left panels of Figure 6, the ratios between the satellite LFs in filaments and in the field become smaller – typically in the range of $1$ to $2$ – compared to those in the galaxy-traced case, which range from $1$ to $\sim 6$. In addition, the ratios are quite similar across different magnitude bins, in contrast to the galaxy-traced case, where environmental differences become more pronounced for brighter central galaxies. When the central galaxy samples are resampled to match their $M_{r,{\rm cen}}$ distributions (middle panels), the environmental differences decrease, similar to what is seen in the galaxy-traced case. Further resampling to match $M_{\rm 200c}$ distributions (right panels) leads to an even larger reduction, with the environmental differences nearly disappearing within the error bars. The average ratios are summarized in Table 2. Compared to the original environmental differences, matching both the $M_{r,{\rm cen}}$ and $M_{\rm 200c}$ distributions results in relative decreases of $95\%$, $96\%$, and $60\%$ in the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins, respectively – comparable to the results from the galaxy-traced case.
Therefore, the difference in satellite abundance in dark matter–traced filaments is less than one third of that found in galaxy-traced filaments, indicating that tracing filaments with dark matter reduces the environmental dependence of satellite LFs by more than $70\%$. However, the relative contributions from differences in magnitude and halo mass distributions remain similar, together accounting for approximately $\sim 60\%$ to $95\%$ of the effect, depending on the central galaxy magnitude bin.
4.3 How do galaxy tracers enhance the satellite abundance in filaments?
The effects of tracer choice on satellite LFs, outlined in Section 4.2, are closely related to the properties of filament-finding algorithms. Algorithms such as DisPerSE  construct the cosmic web based on the spatial distribution of tracer particles. They estimate a density field from the tracer distribution and identify filaments as ridge-like structures connecting local density maxima. As a result, regions with more densely clustered tracers yield higher estimated densities and are more likely to be identified as part of the filamentary network. Consider two halos with similar virial masses residing in comparable large-scale environments – one hosting more satellites (satellite-rich) and the other fewer (satellite-poor), although their actual surrounding dark matter densities may be similar, the satellite-rich halo contributes more to the galaxy-traced density field and is therefore more likely to be classified as residing within a galaxy-traced filament.
To illustrate this intuitively, we construct controlled central galaxy samples – satellite-rich and satellite-poor – based on the filament galaxy samples from both the galaxy-traced and dark matter-traced cases described earlier. We adopt the same three central galaxy magnitude bins used previously: $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20. In each magnitude bin, satellite-rich central galaxies are defined as those hosting at least one satellite with $M_{\rm diff}<3$, while satellite-poor central galaxies host none. Both central galaxy samples are then further resampled to match the $M_{r,{\rm cen}}$ and virial mass $M_{\mathrm{200c}}$ distributions, ensuring a fair comparison.
For the galaxy-traced filaments, the number ratios between satellite-rich and satellite-poor central galaxies are $N_{\rm rich}/N_{\rm poor}=$ 1.28, 1.37, and 1.25 in the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins, respectively. In contrast, these ratios are $N_{\rm rich}/N_{\rm poor}=$ 0.97, 0.96, and 1.05 in the dark matter-traced filaments (see markers in Figure 7). This clearly shows that, compared to the dark matter-traced case, the galaxy-traced filaments tend to pass through central galaxies hosting more satellites. This explains the more pronounced environmental differences in satellite LFs observed for the galaxy-traced case in previous subsections. Additionally, the fact that the ratio $N_{\rm rich}/N_{\rm poor}$ is quite close to $1$ in the dark matter-traced case further suggests that dark matter density is a less biased tracer of the underlying filamentary network.
Note that this result is fairly robust with respect to the DisPerSE parameter choices in filament identification. For example, we have tested various persistence ratio thresholds – which affect the total filament length – in both tracer choices, and found the results to hold across a wide range of parameter values. In Figure 7, we plot the $N_{\rm rich}/N_{\rm poor}$ ratio as a function of the total filament length normalized to the fiducial case, $L_{\rm tot,test}/L_{\rm tot,fiducial}$ (a proxy for the persistence ratio threshold). Increasing the persistence threshold reduces both the number and total length of filaments, as well as the number of central galaxies residing within them. Conversely, lowering the threshold includes more uncertain filaments, increasing the total filament length and altering the $L_{\rm tot,test}/L_{\rm tot,fiducial}$ ratio. Despite the total filament length varying by over an order of magnitude (from $L_{\rm tot,test}/L_{\rm tot,fiducial}\sim 0.1$ to $\sim 3.5$), galaxy-traced filaments consistently show a higher likelihood of including satellite-rich central galaxies. In contrast, dark matter-traced filaments yield comparable counts of satellite-rich and satellite-poor galaxies across all thresholds. This test reinforces that galaxy tracers systematically bias filament classification toward central galaxies with more satellites, amplifying the environmental dependence of satellite abundance.
5 Summary
We utilize the Illustris TNG100-1 hydrodynamical simulations to study the impact of cosmic filaments on satellite galaxy abundance. Filamentary structures are identified using the DisPerSE  algorithm. Our main findings are summarized as follows.
(i) Environmental dependence. When filaments are identified using galaxies as tracers, we find that central galaxies in filaments tend to host more satellite galaxies than their counterparts in the field, in qualitative agreement with SDSS observational results (Figure 2). Specifically, the average ratios between satellite LFs in filaments and those in the field are $3.49$, $2.61$, and $1.90$ in the central galaxy magnitude bins of $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20, respectively (Table 2).
(ii) Impact of magnitude and halo mass distributions. Within each central galaxy magnitude bin, the differences in the $M_{r,{\rm cen}}$ and host halo mass ($M_{\rm 200c}$) contribute significantly to the measured environmental effects. After resampling to match the $M_{r,{\rm cen}}$ distributions in the two environments, the differences in their satellite LFs are reduced by up to $17\%$. Further resampling to match the $M_{\rm 200c}$ distributions decreases the environmental differences by up to $79\%$ (Figures 3 and 4).
(iii) Impact of filament tracer choice. The choice of filament tracers also plays a significant role in the environmental differences in satellite LFs. Compared to the galaxy-traced case, the environmental differences in the filamentary network identified using dark matter tracers are reduced by $\sim 70\%$. However, the relative contributions from differences in magnitude and halo mass distributions remain similar, together accounting for approximately $\sim 60$$95\%$ of the effect on satellite abundance (Figure 6).
(iv) Cause of tracer effects. The enhanced environmental effects in the galaxy-traced case is closely related to the properties of filament-finding algorithms. Compared to the dark matter density tracers, using galaxies as tracers systematically biases filament classification toward central galaxies with more satellites, amplifying the observed environmental dependence of satellite abundance (Figure 7).
In summary, by taking advantage of hydrodynamical simulations, we have explored the impact of halo mass distributions and filament tracer choice on the environmental dependence of satellite LFs. Our results indicate that much of the environmental differences between filament and field satellites are caused by differences in the halo mass distribution. These findings are useful for understanding and interpreting the observational results of satellite galaxy populations.
We acknowledge support from the National Key Program for Science and Technology Research and Development of China (2023YFB3002500), the National Natural Science Foundation of China (NSFC; No. 12588202) and the China Manned Space Project (CMS-CSST-2025-A03). SL acknowledges the support by the NSFC grant (No. 12473015). LW acknowledges support from the National SKA Program of China (No.2022SKA0110201), and the National Key Research and Development Program of China (No.2023YFB3002501).

### v2

@@
 show]yxmeng@bao.ac.cn
-1 Introduction
+show]shliao@nao.cas.cn
+I Introduction
 […]
@@
 The paper is organized as follows.
-In Section 2, we describe the details of the IllustrisTNG simulations, the filament identification algorithm, and the definition of environment.
-In Section 3, we present the impact of filamentary environment on satellite abundance, including comparisons with observations, the impact of magnitude distributions, and the influence of halo mass distributions.
-The effects of tracer choice are investigated in Section 4.
-Finally, we summarize our findings and conclude in Section 5.
-2 Methods
-2.1 Simulation
-Table 1: Number of central galaxies in bins of $r$-band absolute magnitude ($M_{r,{\rm cen}}$) and different cosmic web environments.
-Bins $M_{r,{\rm cen}}$ range Total Knot Galaxy tracers Dark matter tracers
+In Section II, we describe the details of the IllustrisTNG simulations, the filament identification algorithm, and the definition of environment.
+In Section III, we present the impact of filamentary environment on satellite abundance, including comparisons with observations, the impact of magnitude distributions, and the influence of halo mass distributions.
+The effects of tracer choice are investigated in Section IV.
+Finally, we summarize our findings and conclude in Section V.
+II Methods
+II.1 Simulation
+Table 1: Number of central galaxies (and their associated satellites) in bins of $r$-band absolute magnitude ($M_{r,{\rm cen}}$) and in different cosmic web environments.
+Each bin includes all centrals with $M_{r,{\rm cen}}$ within $\pm 0.5$ mag of the bin center.
+The numbers in parentheses give the total number of satellites (with more than 100 stellar particles and located within the $R_{\rm 200c}$ of their host) associated with the centrals in each bin.
+The listed stellar mass ranges correspond to the 5th–95th percentile of all central galaxies, irrespective of environment.
+Note that the centrals in each bin are selected by $M_{r,{\rm cen}}$, so their stellar mass ranges slightly overlap.
+For both tracers, the total number of central galaxies in the combined filament and field environments is the same in each magnitude bin, owing to our environment definition (see Section II.3).
+Bins $M_{\rm star}$ range Total Knot Galaxy tracers Dark matter tracers
 Filament Field Filament Field
-$M_{r,{\rm cen}}\sim-23$ [23.5,22.5)[-23.5,-22.5) 64 39 23 2 24 1
-$M_{r,{\rm cen}}\sim-22$ [22.5,21.5)[-22.5,-21.5) 455 19 235 201 328 108
-$M_{r,{\rm cen}}\sim-21$ [21.5,20.5)[-21.5,-20.5) 1654 47 379 1228 756 851
-$M_{r,{\rm cen}}\sim-20$ [20.5,19.5)[-20.5,-19.5) 2628 104 316 2208 760 1764
+$M_{r,{\rm cen}}\sim-23$ $10^{11.5-12.1}M_{\odot}$ 64 (3448) 39 (2960) 23 (469) 2 (19) 24 (482) 1 (6)
+$M_{r,{\rm cen}}\sim-22$ $10^{10.8-11.6}M_{\odot}$ 455 (3381) 19 (578) 235 (2187) 201 (616) 328 (2318) 108 (485)
+$M_{r,{\rm cen}}\sim-21$ $10^{10.3-11.0}M_{\odot}$ 1654 (2354) 47 (35) 379 (974) 1228 (1345) 756 (1345) 851 (974)
+$M_{r,{\rm cen}}\sim-20$ $10^{9.8-10.5}M_{\odot}$ 2628 (922) 104 (23) 316 (184) 2208 (715) 760 (326) 1764 (573)
 The analysis in this study is based on the IllustrisTNG simulations (F.
@@
 Dolag et al., 2009) algorithms, respectively.
-We define the host halo mass as $M_{\mathrm{200c}}$, the total mass enclosed within the virial radius $R_{\mathrm{200c}}$, within which the mean density is 200 times the cosmic critical density.
+We define the host halo mass as $M_{\rm 200c}$, the total mass enclosed within the virial radius $R_{\rm 200c}$, within which the mean density is 200 times the cosmic critical density.
 The galaxy sample is based on the SUBFIND subhalo catalogue in the TNG100-1 run.
@@
 Nelson et al.
-2019) as well as low-mass subhalos containing fewer than 100 star particles, which corresponds to satellite galaxies with stellar masses of $M_{\mathrm{star}}\sim 10^{8}~M_{\odot}$ or $r$-band magnitudes $M_{r}\sim-14.6$ (with a typical scatter of $\Delta M_{r}=0.6$).
+2019) as well as low-mass subhalos containing fewer than 100 star particles, which corresponds to satellite galaxies with stellar masses of $M_{\rm star}\sim 10^{8}~M_{\odot}$ or $r$-band magnitudes $M_{r}\sim-14.6$ (with a typical scatter of $\Delta M_{r}=0.6$).
 In this study, the stellar mass and $r$-band magnitude of each SUBFIND structure are computed using star particles within a 3D radial aperture of 30 physical kpc, consistent with the SDSS Petrosian aperture.
@@
 Nelson et al.
-(2018), which are based on a model that accounts for dust attenuation due to the distribution of neutral gas and metals (specifically, ‘Model C’ in D.
+(2018), which are based on a model that accounts for internal dust attenuation arising from the distribution of neutral gas and metals within galaxies (specifically, ‘Model C’ in D.
 Nelson et al.
 2018).
+[^1] We have tested and confirmed that our main conclusions are not sensitive to the dust model and remain consistent when using dust-free $M_{r}$.
 We include all central galaxies with magnitudes in the range $23.5\leq M_{r,{\rm cen}}\leq-19.5$, and divide them into four magnitude bins.
 The bright-end limit of $M_{r,{\rm cen}}=-23.5$ is motivated by G15, while the faint-end limit of $M_{r,{\rm cen}}=-19.5$ is chosen to maximize the sample size within the range of reliable numerical resolution.
-The number of central galaxies in each bin is listed in the third column of Table 1.
-2.2 Filament identification
+The third column of Table 1 lists the total number of central galaxies in each bin (with the corresponding satellite counts given in parentheses), while the second column provides the typical stellar mass range of centrals.
+II.2 Filament identification
 Figure 1: Filaments in a $30$ Mpc-thick slice of the TNG100-1 simulation, identified by the DisPerSE algorithm using galaxies with stellar masses $M_{\rm star}\geq 10^{9}~M_{\odot}$ as tracers.
 The grayscale background shows the dark matter density field.
-White circles mark the $3\times R_{\mathrm{200c}}$ vicinities of galaxy clusters with $M_{\mathrm{200c}}\geq 10^{13.5}~M_{\odot}$, which are defined as the knot environment.
+White circles mark the $3\times R_{\rm 200c}$ vicinities of galaxy clusters with $M_{\rm 200c}\geq 10^{13.5}~M_{\odot}$, which are defined as the knot environment.
 Red dots indicate the centers of galaxies residing in knots, which are excluded from our analysis.
@@
 (2024), we select galaxies according to observational limits, balancing the identification of real structures with the suppression of spurious ones.
-Specifically, we trace filaments using galaxies (including both central and satellite galaxies) with stellar masses $M_{\mathrm{star}}\geq 10^{9}~M_{\odot}$, apply smoothing on both the estimated density field and the identified filaments (each smoothed once), and select filaments with a persistence ratio above the threshold of $2\sigma$.
-Note that we have tested a range of persistence ratio thresholds (from $1\sigma$ to $6\sigma$), with and without smoothing procedures applied, and confirmed that the conclusions presented in the following sections are not sensitive to these choices.
+Specifically, we trace filaments using galaxies (including both central and satellite galaxies) with stellar masses $M_{\rm star}\geq 10^{9}~M_{\odot}$, apply smoothing on both the estimated density field and the identified filaments (each smoothed once), and select filaments with a persistence ratio above the threshold of $2\sigma$.
+Note that we have tested a range of persistence ratio thresholds (from $1\sigma$ to $6\sigma$, see Appendix A for details), with and without smoothing procedures applied, and confirmed that the conclusions presented in the following sections are not sensitive to these choices.
 The identified filament spines in a slice of the simulation box are shown as black solid lines in Figure 1.
@@
 These discrepancies arise primarily from the use of a 2D projection and from galaxies being a sparse tracer of the density field.
-The impact of tracers will be examined in Section 4.
-2.3 Definition of environment
+The impact of tracers will be examined in Section IV.
+II.3 Definition of environment
 After identifying the filamentary spines, we further classify central galaxies into ‘filament’ and ‘field’ environments.
 To do this, we first exclude the galaxies residing in clusters (i.e., the ‘knot’ environment), as these environments have distinct effects on galaxy properties compared to filaments and fields.
-Specifically, central galaxies located within $3R_{\mathrm{200c}}$ of clusters with $M_{\mathrm{200c}}\geq 10^{13.5}~M_{\odot}$ are classified as ‘knot’ environment galaxies (i.e., galaxies marked by red dots in Figure 1) and are excluded from our galaxy samples.
+Specifically, central galaxies located within $3R_{\rm 200c}$ of clusters with $M_{\rm 200c}\geq 10^{13.5}~M_{\odot}$ are classified as ‘knot’ environment galaxies (i.e., galaxies marked by red dots in Figure 1) and are excluded from our galaxy samples.
 The numbers of ‘knot’ galaxies in different $M_{r,{\rm cen}}$ bins are listed in the fourth column of Table 1.
@@
 […]
-3 Satellite luminosity functions
-3.1 Environmental dependence
+III Satellite luminosity functions
+III.1 Environmental dependence
 Figure 2: Environmental dependence of satellite LFs.
@@
 $\phi_{\rm sat}(M_{\rm diff})=\frac{\sum_{i=1}^{N_{\rm cen}}N_{{\rm sat},i}(M_{\rm diff})}{N_{\rm cen}\Delta M_{\rm diff}},$ (1)
-where $N_{{\rm sat},i}(M_{\rm diff})$ is the number of satellites associated with the $i$-th central galaxy whose magnitude differences fall within the bin $M_{\rm diff}-\Delta M_{\rm diff}/2,M_{\rm diff}+\Delta M_{\rm diff}/2)$, and $N_{\rm cen}$ is the total number of central galaxies considered.
+where $N_{{\rm sat},i}(M_{\rm diff})$ is the number of satellites associated with the $i$-th central galaxy whose magnitude differences fall within the bin $M_{\rm diff}-\Delta M_{\rm diff}/2,M_{\rm diff}+\Delta M_{\rm diff}/2)$, $\Delta M_{\rm diff}$ denotes the bin width of the magnitude difference, and $N_{\rm cen}$ is the total number of central galaxies considered.
 We also estimate the uncertainty in the satellite LF, shown as shaded regions, based on Poisson errors, i.e.,
@@
 These average ratios are summarized in the third column of Table 2.
+Figure 3: Top: Distributions of central galaxy magnitudes ($M_{r,{\rm cen}}$).
+From left to right, we show the distributions of the original samples, the samples after resampling $M_{r,{\rm cen}}$, and the samples after resampling both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$.
+In each panel, as in other figures in this study, the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins are plotted in red, green, and yellow, respectively.
+Central galaxies in filaments and the field are distinguished using solid and dashed lines, respectively.
+For each $M_{r,{\rm cen}}$ bin, the Kolmogorov–Smirnov (KS) test $p$-value is shown in the panel, quantifying the similarity between the filament and field $M_{r,{\rm cen}}$ distributions.
+Bottom: Same as the top panels, but showing the distributions of the host halo virial masses ($M_{\rm 200c}$) for central galaxies.
+After resampling, the filament and field samples exhibit similar distributions in both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$.
+See the main text for details of the resampling procedure.
 Overall, the environmental difference is quantitatively more pronounced in simulations compared to the G15 observations.
@@
 In the following subsections, we further examine the possible origin of this environmental difference.
-Figure 3: Top: Distributions of central galaxy magnitudes ($M_{r,{\rm cen}}$).
-From left to right, we show the distributions of the original samples, the samples after resampling $M_{r,{\rm cen}}$, and the samples after resampling both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$.
-In each panel, as in other figures in this study, the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins are plotted in red, green, and yellow, respectively.
-Central galaxies in filaments and the field are distinguished using solid and dashed lines, respectively.
-Bottom: Same as the top panels, but showing the distributions of the host halo virial masses ($M_{\rm 200c}$) for central galaxies.
-After resampling, the filament and field samples exhibit similar distributions in both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$.
-See the main text for details of the resampling procedure.
 Figure 4: Impact of magnitude and halo mass distributions on the environmental differences in satellite LFs.
@@
 Filament tracers Bins $\phi_{\rm sat,filament}/\phi_{\rm sat,field}$
-Original Resampling $M_{r,{\rm cen}}$ Resampling $M_{r,{\rm cen}}$ and $M_{\mathrm{200c}}$
+Original Resampling $M_{r,{\rm cen}}$ Resampling $M_{r,{\rm cen}}$ and $M_{\rm 200c}$
 Galaxies $M_{r,{\rm cen}}\sim-22$ $3.49$ $3.07~(17\%\downarrow)$ $1.55~(78\%\downarrow)$
@@
 […]
-3.2 Impact of magnitude distributions
+III.2 Impact of magnitude distributions
 In the previous analysis, we adopted a bin width of $\Delta M_{r,{\rm cen}}=1$ for central galaxies.
@@
 Overall, central galaxies in filaments do show a higher fraction of brighter members compared to those in the field, which can partially explain their higher satellite abundance.
-To further quantify the impact of magnitude distribution on the environmental difference observed in Section 3.1, we resample the central galaxies based on their $M_{r,{\rm cen}}$, ensuring that the resampled samples have similar distributions in both environments.
+To further quantify the impact of magnitude distribution on the environmental difference observed in Section III.1, we resample the central galaxies based on their $M_{r,{\rm cen}}$, ensuring that the resampled samples have similar distributions in both environments.
 Specifically, for each pair of filament and field histograms, we divide them into 20 narrower bins.
 [^2] We have tested other choices for the number of narrow bins (e.g., 10, 40) and confirmed that the results remain consistent.
-A similar test is also performed for the resampling of $M_{\rm 200c}$ (Section 3.3).
+A similar test is also performed for the resampling of $M_{\rm 200c}$ (Section III.3).
 In each narrow bin, central galaxies from the environment with a higher number count are randomly resampled to match the number of galaxies in the other environment.
 The $M_{r,{\rm cen}}$ distributions of the resampled central galaxy samples are shown in the upper-middle panel of Figure 3.
-We can clearly see that, after resampling, the filament and field histograms are closely matched.
+We can clearly see that, after resampling, the filament and field histograms are closely matched, as further confirmed by the Kolmogorov–Smirnov (KS) tests shown in Figure 3 (i.e., a higher $p$-value indicating that the two distributions are statistically indistinguishable).
 We then re-examine the environmental differences in satellite LFs using these resampled central galaxies.
@@
 Compared to the original ratios, the resampled values decrease by less than $17\%$.
-This suggests that the difference in magnitude distributions accounts for only a small part of the observed environmental effect, and that the environmental difference reported in Section 3.1 likely arises from other factors.
-3.3 Impact of halo mass distributions
+This suggests that the difference in magnitude distributions accounts for only a small part of the observed environmental effect, and that the environmental difference reported in Section III.1 likely arises from other factors.
+III.3 Impact of halo mass distributions
 Another factor that can impact the abundance of satellite galaxies is the host halo mass, $M_{\rm 200c}$ (e.g., R.
@@
 This indicates that even after matching the magnitude distributions, the halo mass distributions between filaments and the field remain substantially different, which could contribute to the environmental differences observed in the satellite LFs.
-To quantify the impact of halo mass distributions, we follow the procedure described in Section 3.2.
+To quantify the impact of halo mass distributions, we follow the procedure described in Section III.2.
 Starting from the resampled central galaxy samples with matched $M_{r,{\rm cen}}$ distributions, we further divide the halo mass distributions (shown in the bottom-middle panel of Figure 3) into 20 narrower bins.
@@
 This suggests that the differences in halo mass distributions between filament and field central galaxies make a significant contribution to the environmental differences in satellite abundance observed in the original samples, especially in the brighter central galaxy bins.
-3.4 Discussions
+III.4 Discussions
 As shown in the previous subsections, the environmental differences in satellite abundances are partly driven by variations in both central galaxy magnitude and host halo mass distributions.
@@
 In particular, using galaxies as tracers may enhance the environmental dependence of satellite LFs.
-4 Tracer Effects
+IV Tracer Effects
 Figure 5: Similar to Figure 1, but for filaments identified using dark matter density as tracers.
@@
 Filaments identified by these two tracers are qualitatively similar, both following the underlying filamentary matter distribution, but they differ in detailed structures.
-We investigate these differences and their impact in Section 4.
+We investigate these differences and their impact in Section IV.
 Figure 6: Same as Figure 4, but for the parallel analysis where filaments are identified using dark matter density as the tracer.
@@
 In contrast, dark matter-traced filaments exhibit ratios close to unity (gray dashed horizontal line), indicating no significant preference.
-4.1 Dark matter-traced filaments
+IV.1 Dark matter-traced filaments
 It has been shown that the choice of tracers can affect the detailed properties of filaments identified by DisPerSE (see e.g., C.
@@
 M.
-Bahe & P.
+Bahé & P.
 Jablonka, 2025).
@@
 This discrepancy arises because galaxies, distributed sparsely and discontinuously, are biased tracers of the underlying matter distribution, whereas dark matter more faithfully traces the total mass density field.
-In observations, we are limited to using galaxies as tracers, as was done in Section 3.
+In observations, we are limited to using galaxies as tracers, as was done in Section III.
 In this section, we take advantage of simulations to explore the use of dark matter as tracers and investigate how the tracer choice affects the observed environmental dependence of satellite galaxy abundance.
@@
 M.
-Bahe & P.
+Bahé & P.
 Jablonka (2025), we use a smoothed density field instead of randomly sampled dark matter particles.
@@
 (2023), and to enable a more direct and fair comparison with galaxy-traced filaments, we adopt a higher persistence ratio threshold of $5.1\sigma$ (compared to $2\sigma$ for galaxy tracers), ensuring that the total filament lengths are matched between the two cases (i.e., $L_{\rm tot,fiducial}=7.84\times 10^{3}$ Mpc).
+Alternatively, we have tested selecting the persistence ratio to match the number of galaxies in filaments to that of the galaxy-traced case, and found that this does not affect our main conclusions.
 The resulting dark matter-traced filament spines are shown in Figure 5 as purple solid lines.
@@
 Especially, the dark matter-traced filaments tend to extend more continuously, covering more low-density ridges that galaxy-traced filaments may miss.
-4.2 Tracer effects on satellite luminosity functions
-Once the dark matter-traced filament spines are identified, we follow the same environmental definitions outlined in Section 2.3 and apply the analysis described in Section 3, as done for the galaxy-traced case.
+IV.2 Tracer effects on satellite luminosity functions
+Once the dark matter-traced filament spines are identified, we follow the same environmental definitions outlined in Section II.3 and apply the analysis described in Section III, as done for the galaxy-traced case.
 The numbers of central galaxies classified into filament and field environments across different magnitude bins are given in the two rightmost columns of Table 1.
+We note that the total number of central galaxies in the filament plus field environments is identical to that in the galaxy-traced case owing to our environment definition (see Section II.3).
 Compared to the galaxy-traced case, a larger number of central galaxies are assigned to the filament environment, reflecting the fact that dark matter-traced filaments can capture more low-density ridges and thus include the galaxies surrounding these structures.
@@
 However, the relative contributions from differences in magnitude and halo mass distributions remain similar, together accounting for approximately $\sim 60\%$ to $95\%$ of the effect, depending on the central galaxy magnitude bin.
-4.3 How do galaxy tracers enhance the satellite abundance in filaments?
-The effects of tracer choice on satellite LFs, outlined in Section 4.2, are closely related to the properties of filament-finding algorithms.
+IV.3 How do galaxy tracers enhance the satellite abundance in filaments?
+The effects of tracer choice on satellite LFs, outlined in Section IV.2, are closely related to the properties of filament-finding algorithms.
 Algorithms such as DisPerSE  construct the cosmic web based on the spatial distribution of tracer particles.
@@
 Consider two halos with similar virial masses residing in comparable large-scale environments – one hosting more satellites (satellite-rich) and the other fewer (satellite-poor), although their actual surrounding dark matter densities may be similar, the satellite-rich halo contributes more to the galaxy-traced density field and is therefore more likely to be classified as residing within a galaxy-traced filament.
-To illustrate this intuitively, we construct controlled central galaxy samples – satellite-rich and satellite-poor – based on the filament galaxy samples from both the galaxy-traced and dark matter-traced cases described earlier.
+To illustrate this intuitively, we construct controlled central galaxy samples – satellite-rich and satellite-poor – from the total central galaxy sample after excluding those in the knot environment (see Section II).
 We adopt the same three central galaxy magnitude bins used previously: $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20.
 In each magnitude bin, satellite-rich central galaxies are defined as those hosting at least one satellite with $M_{\rm diff}<3$, while satellite-poor central galaxies host none.
-Both central galaxy samples are then further resampled to match the $M_{r,{\rm cen}}$ and virial mass $M_{\mathrm{200c}}$ distributions, ensuring a fair comparison.
+[^3] We also tested higher $M_{\rm diff}$ thresholds to include fainter satellites (e.g.
+$M_{\rm diff}<6$) and considered defining satellite-rich centrals as hosting at least one or two satellites.
+In all cases, the main conclusions remain consistent.
+We then further resample the satellite-rich and satellite-poor central galaxy samples to match their $M_{r,{\rm cen}}$ and virial mass $M_{\rm 200c}$ distributions, ensuring a fair comparison.
+Finally, we examine the numbers of satellite-rich and satellite-poor central galaxies in this controlled sample that reside in the identified filaments, denoted $N_{\rm rich}$ and $N_{\rm poor}$, respectively.
 For the galaxy-traced filaments, the number ratios between satellite-rich and satellite-poor central galaxies are $N_{\rm rich}/N_{\rm poor}=$ 1.28, 1.37, and 1.25 in the $M_{r,{\rm cen}}\sim-22$, 21-21, and 20-20 bins, respectively.
 In contrast, these ratios are $N_{\rm rich}/N_{\rm poor}=$ 0.97, 0.96, and 1.05 in the dark matter-traced filaments (see markers in Figure 7).
+[^4] To ensure that difference in the magnitude distributions of central galaxies (the combined satellite-rich and satellite-poor populations) between galaxy-traced and dark matter-traced filaments does not bias the results, we performed a supplemental resampling test.
+After matching both samples in $M_{r,{\rm cen}}$ and $M_{\rm 200c}$, the resulting $N_{\rm rich}/N_{\rm poor}$ ratios remained consistent with those reported in the main text.
 This clearly shows that, compared to the dark matter-traced case, the galaxy-traced filaments tend to pass through central galaxies hosting more satellites.
@@
 This test reinforces that galaxy tracers systematically bias filament classification toward central galaxies with more satellites, amplifying the environmental dependence of satellite abundance.
-5 Summary
+V Summary
 We utilize the Illustris TNG100-1 hydrodynamical simulations to study the impact of cosmic filaments on satellite galaxy abundance.
@@
 These findings are useful for understanding and interpreting the observational results of satellite galaxy populations.
-We acknowledge support from the National Key Program for Science and Technology Research and Development of China (2023YFB3002500), the National Natural Science Foundation of China (NSFC; No.
-12588202) and the China Manned Space Project (CMS-CSST-2025-A03).
+We thank the anonymous referee for their constructive comments, which helped improve this manuscript.
+We acknowledge support from the National Natural Science Foundation of China (NSFC) grant (No.
+12588202), the National Key Program for Science and Technology Research and Development of China (2023YFB3002500), and the China Manned Space Project (CMS-CSST-2025-A03).
 SL acknowledges the support by the NSFC grant (No.
@@
 LW acknowledges support from the National SKA Program of China (No.2022SKA0110201), and the National Key Research and Development Program of China (No.2023YFB3002501).
+Appendix A Effects of filament persistence thresholds
+Filament identification with DisPerSE depends on the choice of the persistence ratio, expressed in units of $\sigma$.
+In the main text, we adopt a fiducial threshold of $2\sigma$.
+Higher thresholds select fewer filaments and filter out unreliable or spurious structures, which results in shorter total filament lengths.
+Conversely, lower thresholds include more filaments and lead to longer total lengths.
+As a result, both the number of galaxies classified as in filaments and the corresponding satellite LFs could be affected by the choice of persistence threshold.
+Here we test the robustness of our conclusions using persistence thresholds of $1\sigma$, $4\sigma$, and $6\sigma$.
+The total filament lengths, $L_{\rm tot}$, at $1\sigma$, $4\sigma$, and $6\sigma$ are approximately $1.65$, $0.65$, and $0.26$ times the fiducial $L_{\rm tot,fiducial}$ at $2\sigma$, respectively.
+The persistence ratios for the dark matter–traced filaments are chosen to match the total filament lengths, as described in Section IV.1.
+Figures 8, 9, and 10 show the environmental differences in the satellite LFs for different persistence thresholds.
+Relative to Figures 4 and 6 (the $2\sigma$ results), the satellite LFs and the ratios between filament and field environments remain broadly similar.
+A more noticeable deviation appears at $M_{r,{\rm cen}}\sim-22$ (original sample, red curve) in the $6\sigma$ case, where many filaments become classified as field regions, causing the environmental ratio to drop.
+In all thresholds considered, resampling in $M_{r,{\rm cen}}$ and $M_{\rm 200c}$ suppresses the environmental differences, and using dark matter as the filament tracer further reduces them.
+For a quantitative comparison, we summarize the average ratios in Table 3, analogous to Table 2.
+We find that the average satellite abundance ratio correlates with the persistence threshold: lower persistence (more filaments and longer total length) corresponds to a higher ratio, and vice versa.
+However, the fractional reduction from resampling remains similar across thresholds.
+Resampling in $M_{r,{\rm cen}}$ reduces the ratio by $\sim 0-30\%$ (16–84th percentile), with a median reduction of $\sim 15\%$.
+Resampling in both $M_{r,{\rm cen}}$ and $M_{\rm 200c}$ reduces the ratio by $\sim 50-90\%$ (16–84th percentile), with a median of $\sim 75\%$.
+For fixed total filament length, dark matter–traced filaments consistently show lower satellite abundance ratios than galaxy-traced filaments.
+In summary, varying the persistence ratio from $1\sigma$ to $6\sigma$ changes the total filament length and correspondingly affects the satellite LFs and abundance ratios.
+Nevertheless, the main conclusions regarding environmental effects and tracer dependence remain robust.
+Figure 8: Environmental differences in satellite LFs, analogous to Figures 4 and 6, but using filaments identified with a persistence threshold of $1\sigma$.
+Left: Galaxy-traced filaments without resampling.
+Middle: Galaxy-traced filaments after resampling in $M_{r,{\rm cen}}$ and $M_{\rm 200c}$.
+Right: Dark matter–traced filaments ($4.4\sigma$) after the same resampling.
+The satellite LFs and their filament-to-field ratios are quantitatively different compared to the fiducial $2\sigma$ case, but the behaviors are qualitatively similar, especially the effects of halo mass and tracer choice are consistent with the results discussed in the main text.
+Figure 9: Same as Figure. 8, but for filaments identifies with a higher persistence threshold of $4\sigma$.
+Figure 10: Same as Figure. 8, but for filaments identifies with a higher persistence threshold of $6\sigma$.
+Table 3: Average satellite abundance ratios between filament and field central galaxies, analogous to Table 2, but for filaments identified with different persistence thresholds ($1\sigma$, $4\sigma$, and $6\sigma$).
+Note that the original ratio in the last row does not exceed unity, so the calculation of the fractional reduction is not applicable and is therefore omitted.
+Filament tracers Bins $\phi_{\rm sat,filament}/\phi_{\rm sat,field}$
+Original Resampling $M_{r,{\rm cen}}$ Resampling $M_{r,{\rm cen}}$ and $M_{\rm 200c}$
+Galaxies ($1\sigma$) $M_{r,{\rm cen}}\sim-22$ $3.62$ $3.19~(17\%\downarrow)$ $1.60~(77\%\downarrow)$
+$M_{r,{\rm cen}}\sim-21$ $2.73$ $2.59~(8\%\downarrow)$ $1.44~(74\%\downarrow)$
+$M_{r,{\rm cen}}\sim-20$ $1.89$ $1.89~(0\%\downarrow)$ $1.40~(56\%\downarrow)$
+Dark matter density ($4.4\sigma$) $M_{r,{\rm cen}}\sim-22$ $3.26$ $2.42~(37\%\downarrow)$ $1.26~(89\%\downarrow)$
+[…]
+$M_{r,{\rm cen}}\sim-20$ $1.64$ $1.54~(15\%\downarrow)$ $1.23~(63\%\downarrow)$
+Galaxies ($4\sigma$) $M_{r,{\rm cen}}\sim-22$ $3.10$ $2.76~(16\%\downarrow)$ $1.50~(76\%\downarrow)$
+$M_{r,{\rm cen}}\sim-21$ $2.59$ $2.46~(9\%\downarrow)$ $1.29~(82\%\downarrow)$
+$M_{r,{\rm cen}}\sim-20$ $1.87$ $1.99~(14\%\uparrow)$ $1.44~(50\%\downarrow)$
+Dark matter density ($5.6\sigma$) $M_{r,{\rm cen}}\sim-22$ $1.61$ $1.35~(42\%\downarrow)$ $1.09~(85\%\downarrow)$
+$M_{r,{\rm cen}}\sim-21$ $1.36$ $1.29~(19\%\downarrow)$ $1.01~(98\%\downarrow)$
+[…]
+Galaxies ($6\sigma$) $M_{r,{\rm cen}}\sim-22$ $1.86$ $1.47~(46\%\downarrow)$ $1.22~(75\%\downarrow)$
+$M_{r,{\rm cen}}\sim-21$ $2.65$ $2.43~(14\%\downarrow)$ $1.32~(81\%\downarrow)$
+$M_{r,{\rm cen}}\sim-20$ $2.20$ $2.47~(22\%\uparrow)$ $1.64~(46\%\downarrow)$
+Dark matter density ($6.5\sigma$) $M_{r,{\rm cen}}\sim-22$ $1.37$ $1.28~(26\%\downarrow)$ $1.00~(100\%\downarrow)$
+$M_{r,{\rm cen}}\sim-21$ $1.33$ $1.30~(9\%\downarrow)$ $1.02~(95\%\downarrow)$
+$M_{r,{\rm cen}}\sim-20$ $0.89$ 0.89()0.89~(-) 1.08()1.08~(-)